OpenStudy (anonymous):

Hitomi has a bag with four chips in it numbered 1 to 4. She randomly draws a chip and replaces it three times. What is the probability that Hitomi will draw a 1 all three times?

OpenStudy (anonymous):

take all of the probabilities and multiply them together.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

will it be improper XD

OpenStudy (anonymous):

Well you should put it in fraction form, not decimal.

OpenStudy (anonymous):

thats what i'm talking about i didn't know there were improper decimals...

OpenStudy (anonymous):

is it an improper fraction?

OpenStudy (anonymous):

1/4*1/4*1.4 will not be an improper fraction.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3/4?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }*\frac{ 1 }{ 4 }*\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

i know but is that the answer....( i really suck at math)

OpenStudy (anonymous):

Well we are multiplying, not adding. So no it's not the answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here, \[\frac{ 1 }{ 4 }*\frac{ 1 }{ 4 }*\frac{ 1 }{ 4 }=\frac{ 1*1*1 }{ 4*4*4 }\]

OpenStudy (anonymous):

1/64

OpenStudy (anonymous):

And that is your answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem :)